# Microscopic Analysis on LLM players via Social Deduction Game

### v1

@@ Abstract
Recent studies have begun developing autonomous game players for social deduction games using large language models (LLMs). When building LLM players, fine-grained evaluations are crucial for addressing weaknesses in game-playing abilities. However, existing studies have often overlooked such assessments. Specifically, we point out two issues with the evaluation methods employed. First, game-playing abilities have typically been assessed through game-level outcomes rather than specific event-level skills; Second, error analyses have lacked structured methodologies. To address these issues, we propose an approach utilizing a variant of the SpyFall game, named SpyGame. We conducted an experiment with four LLMs, analyzing their gameplay behavior in SpyGame both quantitatively and qualitatively. For the quantitative analysis, we introduced eight metrics to resolve the first issue, revealing that these metrics are more effective than existing ones for evaluating the two critical skills: intent identification and camouflage. In the qualitative analysis, we performed thematic analysis to resolve the second issue. This analysis identifies four major categories that affect gameplay of LLMs. Additionally, we demonstrate how these categories complement and support the findings from the quantitative analysis.
@@ I Introduction
Conversation plays a pivotal role in social deduction games (SDGs), such as Among Us and Goose Goose Duck. In these games, players are divided into multiple adversarial groups and interact by executing actions through natural language sentences. Unlike other games, these interactions create tricky conversations where explicit statements often contradict the implicit strategies of the corresponding speaker. For example, in Among Us, imposters explicitly and repeatedly claim to be members of the crew. However, this assertion is part of their strategy to conceal their true identity. Consequently, crew members must uncover these hidden intentions to win. As this example illustrates, SDG players must possess the ability to navigate and manage such tricky communications.
Although such abilities are essential, traditional game agents may struggle to handle the tricky conversations occurring in SDGs. Recent studies have adopted large language models (LLMs) because LLMs are capable of generating fluent dialogues and strategic plans. Researchers believe that these capabilities could make LLMs suitable for supporting SDGs. However, most studies have primarily focused on improving gameplay performance rather than investigating whether LLMs can genuinely support these intricate conversations. Specifically, existing studies have two key issues when evaluating SDG-playing agents: (1) Macroscopic quantitative evaluation and (2) Non-systematic qualitative analysis.
@@
First, the issue of macroscopic evaluation refers to the tendency of prior studies to evaluate various game-playing abilities using broad game-level outcomes instead of focusing on specific event-level skills. Existing research often relies on simple summative metrics, such as winning rates, which fail to account for the events that occur during gameplay[1, 2]. While these studies acknowledge the importance of certain skills for success in SDGs, they may overlook the need to measure these skills separately with such simple summative metrics. To scrutinize a specific skill, we suggest using microscopic analysis augmented with several specialized metrics. By performing a multifaceted formative analysis leveraging these metrics, one can reveal the true conversational abilities of LLMs in SDG play. In this study, we propose eight microscopic metrics to evaluate two essential skills: (1) Intent identification, the ability to combine explicit clues to uncover hidden information, and (2) Camouflage, which is the ability to disguise oneself from being detected as an opponent.
@@
Second, the non-systematic evaluation issue indicates that previous studies lacked structured methods for error analyses. While certain studies identified abnormal patterns through simple qualitative analysis, the generalizability of their results was limited by their analytical procedures. For instance, some studies employed LLMs to analyze data without incorporating human-annotation procedures, while others did not adequately discuss their qualitative analysis methods for deriving generalizable abnormal patterns. To enhance the generalizability of qualitative analysis, we employ a systematic approach known as thematic analysis[3]. Moreover, we aim to draw meaningful insights by correlating the resulting categories with quantitative findings.
@@
We conduct two studies to address these issues. In Study 1 (Section IV), we introduce microscopic metrics and discuss why macroscopic metrics are insufficient for analyzing specific skills. In Study 2 (Section V), we conduct a thematic analysis to derive categories and link these findings to the results of Study 1. In both studies, we let four LLMs play SpyGame, a variant of SpyFall with simplified actions (Section III-A). Consequently, this study makes the following contributions:
  • We demonstrate effectiveness of using microscopic event-level metrics to analyze performance of LLMs in SDGs. Specifically, we show that eight microscopic metrics offer a more detailed assessment than conventional macroscopic metrics in evaluating gameplay skills in SpyGame.
  • We conduct a qualitative analysis to identify categories of abnormal reasoning patterns in LLMs. Through thematic analysis involving multiple LLMs, we discover four main categories and five subcategories, including exposure, role ambiguity, memory distortion, and dissociation.
  • Finally, we validate our quantitative findings by correlating them with the results of the qualitative analysis.
@@ II Related Works
This section reviews the existing literature in two ways. First, we provide a brief overview of LLMs and the historical attempts at developing LLM-based gameplay agents. Then, we point out their two issues on evaluating their agents: macroscopic evaluation and non-systematic analysis.
II-A Existing LLM-based Game Players
Before discussing LLM-based agents in SDGs, we first explain LLMs. An LLM is a language model trained on enormous corpora, excelling in various downstream tasks without requiring task-specific fine-tuning. Researchers in artificial intelligence have reported that LLMs exhibit human-like skills across diverse tasks, including question answering, communication, and planning [4, 5, 6]. For instance, LLMs have been employed to create chatbots that engage in natural, fluent conversations with humans[7]. Therefore, LLMs have been widely adopted in tasks involving natural conversations, such as humor, empathy, and decision-making[8, 9].
Recently, researchers have employed LLMs to design conversational agents for games. Leveraging their natural conversations, studies have reported that LLMs can play simple conversational games at a human level. For example, Brookins and Debacker[10] demonstrated that GPT-3.5 could play the Dictator game[11] and the Prisoner’s Dilemma game[12], displaying preferences for fairness and cooperation that often exceeded those observed in human players.
After confirming the gameplay ability of LLMs in simple conversational games, researchers have turned their attention to a more complex family of games: SDGs. SDGs require a higher level of communication skills like camouflage, which are not essential for playing simpler games. Therefore, researchers have developed autonomous player agents for SDGs, such as Avalon and Werewolf. For instance, Shi et al.[13] designed an LLM-based agent to play Avalon. To build working agents, they incorporated a reflection mechanism[14], which complemented the memory of LLMs with a human-inspired memory mechanism. Similarly, ReCon[15] offers additional hints to assist LLM players in choosing the best action during decision-making. Xu et al.[16] developed an LLM-based Werewolf player using reinforcement learning. They employed a success condition for specific actions as a reward signal to guide the learning procedure. However, these approaches have two key issues when evaluating the performance of their agents, which are discussed in the following section.
II-B Issues in Evaluating LLM Players
A fine-grained evaluation can point out weaknesses of an agent in the gameplay, which is crucial for designing effective LLM players. However, prior studies have typically employed more coarse-grained approaches. Therefore, these studies encounter two main issues in their evaluations: macroscopic quantitative evaluation and non-systematic qualitative analysis.
First, previous studies have leveraged macroscopic metrics for quantitative evaluation. Some researchers have adopted quantitative metrics, whose calculations are solely based on the results of entire game. For example, Xu et al.[2] evaluated LLM-based Werewolf players utilizing the winning rate, total game duration, and total number of camouflage behaviors. Similarly, Light et al.[17] employed winning rates and action success rates to assess gameplay, despite defining several specific gameplay skills, including deductive reasoning, coordination, collaboration, and deception. While these studies successfully measured the overall gameplay ability of an agent, their discussions remained superficial because the evaluation results did not support the existence of specific skills. Specifically, the winning rate (WR) does not provide insights into skills like deception or reasoning skills, as these skills all contribute to victory. Therefore, microscopic metrics are necessary to evaluate specific skills and conduct a multifaceted formative assessment of the gameplay of an LLM agent.
Second, prior studies have employed non-systematic qualitative analyses. Since macroscopic quantitative analysis alone cannot provide sufficient evidence for specific skills, some researchers have adopted qualitative error analysis to assess whether their agents possess certain skills[1, 17, 18]. For instance, Qiao et al.[1] provided a few samples for error analysis without presenting information on frequency of errors. Similarly, Kim and Kim[18] qualitatively analyzed errors made by LLMs in SpyFall. However, these errors are more related to generation error (e.g., contextual or format errors) than to specific gameplay skills. Though these approaches may provide insights into erroneous behaviors in LLM players, the generalizability of the results is questionable since the error may represent a specific but infrequent case. Also, other researchers have adopted language models to conduct quantitative analyses[19, 15]. For instance, Lan et al.[19] used ChatGPT[20] to assess agent gameplay across five aspects, including leadership and persuasion. While this approach can reduce the need for human labor in qualitative analyses, the reliability of these results is questionable due to potential issues with LLMs have reliability issues (e.g., hallucination) in generating answers. Hence, to obtain a comprehensive understanding of the errors made by LLM players, a systematic qualitative analysis of LLM gameplay is necessary.
In summary, the existing literature has two main issues. To address the first issue of macroscopic quantitative measures, we need microscopic measures that can investigate specific skills rather than evaluating overall gameplay ability. To tackle the second issue of non-systematic qualitative analysis, we need to adopt a systematic method for analyzing gameplay logs. In the following sections, especially in Studies 1 and 2, we demonstrate a method for conducting such required analyses using the SDG named SpyGame.
In summary, existing literature has two issues. To address the first issue of macroscopic quantitative measures, we need microscopic measures which can investigate a specific skills instead of whole gameplay ability. Also, to address the second issue of non-systematic qualitative analysis, we need to adopt a systematic way for analyzing gameplay logs. In the following sections, especially in Study 1 and Study 2, we will demonstrate a way to conduct such required analyses, using a SDG called SpyGame.
@@ III Experiment
This section details the process of collecting gameplay logs from LLMs for both quantitative and qualitative analyses. To evaluate performance of vanilla LLM in SDG play, we let LLMs play an SDG through natural language interactions. Section III-A provides an overview of SpyGame, which is based on SpyFall. To collect the gameplay logs of SpyGame, we employed four LLMs, as detailed in Section III-B. The procedure for collecting logs is outlined in Section III-C.
@@ III-A SpyGame
SpyGame is a modified version of SpyFall played by seven players. We altered Spyfall for two key reasons: First, LLMs may have been exposed to the rules and details of Spyfall during their pre-training phase. This exposure can cause LLMs to embed game-related knowledge from pretraining data in their parameters. Therefore, it is likely that these LLMs use this parametric knowledge as a gameplay hint rather than leveraging their inherent abilities. To mitigate this, we renamed the game to SpyGame.
Second, setting a time budget may influence the reproducibility of our analysis. Various factors in the experimental setup (e.g., the hardware used, network status, and model size) can impact inference time. Thus, to enhance reproducibility, we replaced the total time budget of a game with the total number of “turns” within a game.
@@ III-A1 Team formation
Similar to SpyFall, SpyGame features two teams: six citizen players and one spy player. The citizens share a common location, called a hidden location, with each citizen plays a specific role in that location. However, the spy is not given information about the hidden location or the roles of the citizens. In addition, the citizens are unaware of the identity of the spy.
Players must uncover hidden information to win the game. In each turn, one player asks another player a question about the hidden location, and the questioned player must answer. Through these conversations, citizens try to identify the spy. Also, citizens should need to imply their knowledge of the hidden location without revealing too much detail, to avoid being suspected as the spy. Meanwhile, the spy attempts to determine the hidden location by pretending to be a citizen, gathering clues from the utterances of others. The detailed game rules, location-role sets, and prompts used for LLMs are provided in [anonymized for review].
@@ III-A2 Game Flow
Each turn has a designated leader. The turn leader starts the turn by choosing one of two actions: questioning or accusation. If the leader selects the “question” action, they ask another player a question related to the hidden location. The questioned player must then respond based on their assigned role. Alternatively, if the leader opts for the “accusation” action, they pick the player whom the leader most suspects to be the spy and initiate a “Day Vote.” The other players, excluding the accuser and the accused, then vote on whether they agree with the accusation. If all players agree, the game ends immediately. After completing either the “question” or “accusation” action, the questioned or accused player becomes the leader for the next turn.
In every turn after the first, the spy player can publicly reveal their identity and guess the hidden location. Before each turn begins, the system asks the spy player if they are ready to reveal their identity. The game ends immediately when the spy player decides to reveal their identity. Note that in our implementation, instead of directly asking for the decision of the spy, we ask LLMs how certain they are about their guess.
Despite an accusation or an action of the spy, the game may continue until the final turn: Turn 9 in our implementation. At this point, the system initiates the “Final Vote.” All players secretly vote for the player they most suspect of being the spy. Following the vote, the game ends.
At the end of a game, the spy wins if any of the following conditions is met; otherwise, the citizens win:
  • A citizen is eliminated by a Day Vote.
  • The spy publicly reveals their identity and correctly guesses the hidden location.
  • The Final Vote results in multiple spy candidates, with more than one player receiving the most votes.
  • The Final Vote identifies only one spy candidate who is actually a citizen.
III-B Tested Models
We selected four LLMs for this study: GPT-4[4], GPT-3.5-turbo (Chat-GPT), Gemini-pro (1.0)[5], and LlaMA2-70b-chat[6]. We considered three criteria: model size, reasoning performance, and the existence of publicly-available API. Specifically, we selected models with a parameter size over 70 billion and a reasoning performance exceeding 70% accuracy[21] to ensure reasoning ability required for SDG play. Previous studies have reported a general performance trend among these four models: GPT-4 leads in performance, followed by Gemini, GPT-3.5, and LlaMA2.
We employed the following input and output procedures to analyze the vanilla performance of the four LLMs: First, as an input prompt, we provided only the minimal information necessary to play SpyGame. Specifically, the LLMs received inputs, including game rules, role descriptions, conversation history, and instructions for generating the next action. Unlike previous studies, we did not provide any additional information (such as hints) that could affect gameplay. Second, for output generation, we used a two-stage process involving free-form generation and JSON formatting. We requested LLMs to decide the next action in free form with their strategy. Then we instructed GPT-4 to rewrite the free-form output into a structured JSON format, which is compatible with our environment. By adopting these procedures, we attempted to minimize unintended effects stemming from additional information.
III-C Data Collection
To assess the gameplay of the spy players, we categorized citizens into two types: strong and weak. Since the game involves two adversarial teams, the performance of the spy can be influenced by the gameplay of the opposing citizens. If the citizen players are strong enough to quickly identify the spy, the spy may not have enough turns to demonstrate their abilities. Thus, we conducted separate experiments with weak citizens versus a spy and strong citizens versus a spy. Here, we used LLaMA2-70B-chat as the weak citizens and GPT-4 as the strong citizens, based on results from other reasoning tasks[21, 22, 4] and our preliminary experiments.

### v2

@@ Abstract
Recent studies have investigated whether large language models (LLMs) can support obscure communication that requires specialized skills, such as inferring subtext or doublespeak. To conduct the investigation, researchers have used social deduction games (SDGs) as their experimental environment, in which players conceal and infer specific information. However, prior work has often overlooked how LLMs should be evaluated in such settings. Specifically, we point out two issues with the evaluation methods they employed. First, metrics used in prior studies are coarse-grained as they are based on overall game outcomes that often fail to capture event-level behaviors; Second, error analyses have lacked structured methodologies capable of producing insights that meaningfully support evaluation outcomes. To address these issues, we propose a macroscopic and systematic approach to the investigation. Specifically, we introduce seven fine-grained metrics that resolve the first issue. To tackle the second issue, we conducted a thematic analysis and identified four major reasoning failures that undermine LLMs’ performance in obscured communication.
@@ I Introduction
Large language models (LLMs) have shown that they can successfully conduct typical informative communication with humans by leveraging their natural language understanding abilities [1, 2, 3]. Inspired by these reports, there is a growing interest in adapting LLMs for obscured communication as a next step [4]. Here, the obscured communication indicates that the intended meanings of messages present in the communication are obscured or made unclear to understand, usually by controlling the spread of information. For example, in the real world, an imaginary role-play requires obscure communication to boost immersion in such a role-play. Among the required skills, doublespeak is one of the representative skills that enable a participant to achieve an advantageous position in such a play. So, since LLMs with doublespeak ability can support the play, researchers have recently focused on investigating such abilities. [5]
Meanwhile, investigating whether LLMs can conduct obscure communication requires a specialized environment. As ordinary LLM use cases focus on providing information, obscured communication cannot be tested with such ordinary cases. To set a more suitable environment for the examination, researchers have used social deduction games (SDGs) because SDGs force players to use doublespeak to achieve their goals. Specifically, they let LLMs play an SDG and collected the game logs, including LLMs’ decisions and conversations. After, they devised several metrics using particular game actions for evaluating LLMs’ ability of doublespeak. [6] Also, they conducted qualitative analyses to discover erroneous actions or utterances. For example, [7] examined whether LLMs can support obscured communication by letting them play a famous SDG, Werewolf. However, to thoroughly examine LLMs’ doublespeak abilities, these attempts have two issues: (1) Macroscopic quantitative evaluation and (2) Non-systematic qualitative analysis.
@@
First, the issue of macroscopic evaluation refers to the tendency of prior works to use broad game-level outcomes as their metrics instead of specific event-level actions. Existing research often relies on simple metrics, such as win rates, which fail to account for the events that occur during gameplay [6, 7]. While these studies acknowledge the importance of certain skills for supporting obscured communication, they may overlook the need to measure these skills separately with such granular metrics. To scrutinize a specific skill, we suggest using microscopic analysis augmented with several fine-grained metrics. Specifically, we propose seven metrics to evaluate two essential skills regarding doublespeak: (1) Subtext inference, the deductive ability to identify doublespeak of others, and (2) Deceptive control, which is the ability to control information to make others confused.
@@
Second, the non-systematic evaluation issue indicates that previous studies lacked structured methods for error analysis. While certain studies identified abnormal patterns through simple qualitative analysis, the reproducibility of their results was limited by their analytical procedures. For instance, some studies employed LLMs to analyze data without incorporating human-annotation procedures [8, 9], while others did not adequately discuss their qualitative analysis methods for defining abnormal patterns. Also, they simply reported the occurrence tendency of the patterns rather than discussing how they affected their evaluation results [10]. We believe that conducting qualitative analysis in a systematic way provides both reproducibility and interoperability of experimental results. Therefore, we adopt a systematic approach known as thematic analysis [11] to identify recurring reasoning errors that hinder LLMs’ doublespeak. By correlating these qualitative patterns with our quantitative metrics, we uncover how specific reasoning failures impact LLMs’ performance in subtext inference and deceptive control.
@@
We conducted two studies to address these issues. In Study 1 (Section IV), we perform a quantitative analysis using our fine-grained metrics. The result reveals how coarse-grained metrics have overlooked finer aspects of doublespeak. In Study 2 (Section V), we conduct a thematic analysis on collected logs. Based on the result, we deduce several categories and link these findings to the results of Study 1. In both studies, we let four LLMs play SpyGame, a variant of simple SDG called SpyFall (Section III-A). Consequently, this study makes the following contributions:
  • We demonstrate the effectiveness of using microscopic quantitative analysis to assess the doublespeak of LLMs. Specifically, using seven fine-grained metrics, we show possibility of studying doublespeak in LLMs’ SDG play.
  • We identify four categories of abnormal reasoning patterns that harm obscured communication. Also, by combining results from the two studies, we discover a correlation between the results of the two analyses.
@@ II Related Works
This section reviews the existing literature in two ways. First, we outline how recent research has shifted its focus from evaluating LLMs on informative communication tasks to more complex settings that require obscured communication. Then, we point out two issues posed in their assessment: macroscopic evaluation and non-systematic analysis.
II-A Adapting LLMs for Obscured Communication
LLMs have shown human-level performance in various informative communications. For example, researchers widely tested LLMs’ ability to question-answering or summarization [12, 13]. To accurately and reliably provide answers, LLMs for informative communication attempt to leverage true knowledge and suppress false statements during the generation process [14, 15]. So, researchers attempt to build an honest and harmless assistant for informative communication.
Noting that real-world communication includes both informative and obscured communication, researchers have recently turned their interests to obscured communication. The two types of communication differ in how each participant views their opponents. In informative communication, participants trust their opponents and try to establish a cooperative relationship. Meanwhile, participants lack trust in their opponents in obscured communication; thus, they seldom provide valuable information to others and often form adversarial relationships. So, participants require different abilities to achieve their goals in such an adversarial situation; skills like doublespeak are needed. Specifically, participants should manipulate how they present the information they have or capture subtext from the opponent’s utterances.
In developing negotiation or debate situations, the ability to discern the opponent’s intention in obscured communication is crucial, as subtext is often used in such communication. For example, in the context of applying LLMs to negotiation, prior research has investigated how to measure the ability of an LLM to infer the preferences of the opponent by identifying hidden intentions and meanings in the conversation [16]. Another example research applied LLMs to debates in order to identify and leverage weaknesses in the opponent’s arguments. Here, LLMs’ ability to discern intentions plays a crucial role, helping to accurately recognize hidden intentions or logical weaknesses in the opponent’s claims [17].
As obscured communication can be naturally seen in games, researchers have recently evaluated relevant skills using social deduction games (SDGs). As SDGs force players to utilize obscured communication to win the game, an SDG can be a suitable testbed for assessing or quantifying such communication. So, researchers aimed to demonstrate how well their agents can do obscured communication by letting them play SDGs. For example, Wang et al. [9] proposed an LLM agent capable of playing the Avalon game, introducing mechanisms such as dual contemplation to control conversational information. Similarly, Shi et al. [18] proposed another LLM agent that can play Avalon. Other than these zero-shot attempts, researchers used training methods to improve their agents. For instance, Xu et al. [19] trained an LLM agent applying reinforcement learning to play the Werewolf game. Collectively, these works demonstrate that SDGs offer a structured and adversarial context in which LLMs’ handling of obscured communication can be rigorously assessed.
II-B Attempts to assess LLMs in obscured communication
To enhance LLMs’ ability to support obscured communication, it is essential to identify the specific skill areas that require improvement. Therefore, most researchers have conducted evaluations in both quantitative and qualitative ways to reveal the current status of LLMs in obscured communications. For example, Xu et al. [7] and Light et al. [20] proposed automated methods for quantitatively evaluating their LLM-based players who play Werewolf or Avalon games. Specifically, they adopted particular game conditions, such as win rate or action success rate, that occur during their environmental games as their evaluation metrics.
However, we suspect that such metrics are too coarse-grained to adequately capture intermediate behaviors during gameplay. Specifically, the win rate is only available post hoc, limiting its utility for formative assessment. Additionally, the action success rate may fail to account for the multi-player dynamics of SDGs in which performance involves not only task execution but also managing others’ perceptions, such as avoiding suspicion. In other words, achieving a higher win rate or action success rate in an SDG does not entail that the model possesses superior competence across all sub-skills involved in obscured communication.
Other researchers proposed more granular metrics, such as logicality or social behaviors, that can be computed by scoring the reasoning or statements made by LLMs during an Avalon game[9, 8]. However, such metrics proposed in those studies rely on an auxiliary LLM to qualitatively assess reasonings or statements, rather than being grounded in objective in-game events. Therefore, fine-grained quantitative metrics grounded in objective in-game events are essential for evaluating LLMs’ abilities in multi-agent obscured communication, such as subtext inference and deceptive control.
While quantitative evaluation offers valuable insights into measurable performance, it is often insufficient to fully capture the underlying reasoning errors or abnormal behaviors exhibited by LLMs. To address this limitation, researchers have conducted qualitative analyses either as a complement to quantitative evaluation or as a standalone method. Kim [10] presented the frequency of misbehavior of LLMs while playing SpyFall to support their assessment of LLMs in the game. Quia et al. [6] presented some particular cases of LLMs to support their assessment results driven by their evaluation framework using communication games.
However, the methods used in their qualitative analyses were less structured or not related to quantitative results. For example, the categories adopted in [10] related to the generation procedure of LLMs, such as formatting errors, rather than LLMs’ ability to support their obscured environment. To enhance the utility of qualitative analysis, it should serve not only to validate the outcomes of quantitative evaluations but also to provide insights into the underlying causes of results.
In summary, the existing literature has limitations in its quantitative and qualitative analysis methods. To discover whether LLMs can support obscured communication, quantitative analysis with fine-grained metrics and validation of qualitative analysis are necessary.
@@ III Experiment
This section details the process of collecting gameplay logs from LLMs for both quantitative and qualitative analyses. Similar to prior works, we collected gameplay logs of LLMs by letting play SpyGame, which is suitable for observing obscured communication. Section III-A provides an overview of SpyGame, which is based on Spyfall. We employed four LLMs to collect the gameplay logs of SpyGame. Section III-B details the tested models and procedure of data collection.
@@ III-A SpyGame
Figure 1: SpyGame
SpyGame is a variation of Spyfall played by seven players [^1] https://github.com/elu-lab/spygame. To better assess LLMs’ capacity for obscured communication, we modified the original Spyfall in two ways. First, we renamed the game to SpyGame, rather than using Spyfall. LLMs may have been exposed to the rules and details of Spyfall during their pre-training phase. This exposure can cause LLMs to embed game-related knowledge, such as rules or hints, from pretraining data in their parameters. Therefore, it is likely that these LLMs exploit memorized game-specific knowledge from pretraining, rather than demonstrating reasoning-based communicative skills during gameplay. So, to avoid such pretrained knowledge, we changed the name of the game to SpyGame.
Second, we replaced the total time budget of a game with the total number of “turns” within a game. Setting a time budget may influence the reproducibility of our analysis. Various factors in the experimental setup (e.g., the hardware, network status, and model size) can impact inference time. Thus, to enhance reproducibility, we used the number of turns.
@@ III-A1 Team formation
Similar to Spyfall, SpyGame requires two teams: six citizens and one spy. The citizens share a common location, called a hidden location, with each citizen playing a specific character in the location. For example, let us assume that the location is a school. Then, the citizens are assigned to distinct characters, such as a student or a principal. Meanwhile, the spy does not have any information about the location or characters.
Both teams must manage what information they reveal and conceal throughout the game. Citizens should protect the information about the location, and the spy should pretend to be a citizen. So, to win the game, players must uncover hidden information. Simultaneously, all players should avoid being suspected of being a spy. Each citizen’s utterance should reveal just enough to signal alignment with other citizens, while minimizing the risk of exposing the hidden location to the spy. Meanwhile, the spy should gather subtexts from the conversation and deduce the hidden location.
This indicates that the abilities required of both the citizen and the spy are centered around concealing and revealing specific information. Meanwhile, because SpyGame is adversarial in nature, a substantial imbalance in competence between the evaluated player and their opponents can compromise the validity of the evaluation. Moreover, when evaluating a citizen player, the influence of fellow citizens can confound the measurement, making it more difficult to isolate the individual player’s ability. For these reasons, we restrict our evaluation to the spy, who operates somewhat independently. To assess the performance of spy under varying team conditions, we constructed two scenarios: one with strong citizens and one with weak citizens.
@@ III-A2 Game Flow
SpyGame progresses over multiple turns, with each turn consisting of a specific game event. Figure 1 illustrates the four key events: (a) questioning and answering, (b) accusation and day vote, (c) guessing, and (d) final vote. In each turn, a designated leader initiates the turn by choosing one of two actions: questioning or accusation. If the leader chooses to question, they ask another player about the hidden location (Figure 1a). Then, the questioned player responds based on their assigned character.
If the leader opts for an accusation, they point out the player they suspect to be the spy, triggering a Day Vote (Figure 1b). All players except the accuser and the accused participate in the vote. If all players agree, the game ends immediately. After completing either questioning or accusation, the questioned or accused player becomes the leader for the next turn.
From the second turn, the spy player guesses the hidden location by herself (Figure 1c). For each turn, she tells her guess to the system secretly and provides a number indicating how certain she is. When the certainty is high enough [^2] In our implementation, we use 9 as the threshold out of 10., the system announces her guess and ends the game. Otherwise, the game continues.
To avoid endless play, we set a final turn [^3] In our implementation, we set the last turn as 9.. When players finish the final turn, the “Final Vote” begins (Figure 1d). All players secretly vote for the most suspicious player in their minds. At the end of a game, the spy wins when one of the following conditions is satisfied. Otherwise, the citizens win.
  • All players agreed to eliminate a citizen in a Day Vote.
  • After the Final Vote, a citizen has the highest vote without a tie.
  • The spy’s guess is correct when the system announces the spy’s guess.
III-B Models and Data
We selected four LLMs for our experiment: GPT-4[1], GPT-3.5-turbo (ChatGPT), Gemini-pro 1.0[3], and Llama2-70b-chat[2]. As obscured communication requires reasoning on subtexts, we selected large enough models with sufficient reasoning ability on simple tasks. Specifically, we selected models with a parameter size over 70 billion and a reasoning performance exceeding 70% accuracy[21] in 2023. LLMs provide their actions or utterances with their reasoning in natural language. So, we employed GPT-4 to sanitize their outputs into a machine-readable format.